Ticket Number: HEA-NOC/20090331-2 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20090331 12:05 UTC+1 Problem Start: 20090401 08:15 UTC+1
Ticket Update: 20090401 13:57 UTC+1 Problem End: 20090401 09:00 UTC+1
Site/line: HEA-NOC, NCIRL
A new OADM has been installed in the HEAnet GD5 office to provide a
wavelength to the testlab. To add this OADM, there will be a brief outage
on the UPC fibre ring (TCD, GD5, NCI) while patching is performed.
NCI: loss of resilience (NCI-TCD fibre will be unaffected)
GD5: outage to gd5-sw1 for ~5mins. Loss of resilience to HEAnet office IP
connectivity (cpe1-heanet has a Colt fibre link to kp-sw1). Office phone
20090331 11:37 CS Scheduling work.
20090401 09:11 CS Work complete, will switch HEAnet office link from
backup to primary at 09.25. This will cause a brief outage.
20090401 13:52 CS HEAnet office link switched back at 09.25. Asymmetric v6
routing for the office was discovered, but is unrelated to this work and is
Have compared RX power levels on all SFPs using the fibre ring to their
levels before work - no problems. Will keep ticket open until tomorrow for
This ticket can be monitored at http://www.hea.net/tickets/20090331-2
HEAnet Limited. Registered in Ireland, No. 275301.